# Threats, Attacks, and Defenses in Machine Unlearning: A Survey

### v3

@@ IV-B Value Alignment
In these efforts, unlearning is applied in various innovative ways to align AI operations with ethical and legal standards. For instance, attribute unlearning [90] is used to remove sensitive attributes for privacy compliance. Similarly, unlearning with corresponding detection methods are applied to LLMs to remove behaviors and content that are deemed inappropriate or should be avoided, effectively eliminating illegal, poor quality, copyrighted, or harmful content [39, 91, 92, 93, 94, 95, 96, 97]. Differently, [98] employs unlearning to identify the training datasets most influential in generating harmful content. By unlearning each dataset through gradient ascent, the impact on the model’s harmful content generation can be assessed post-unlearning.
V Evaluating Unlearning through Attacks
@@
V-A Audit of Privacy Leakage
As outlined in Section III-A, machine unlearning systems are susceptible to threats of information leakage. Attackers can exploit discrepancies between the trained and unlearned models, or inherent vulnerabilities within the unlearning system, to initiate inference attacks [29, 51, 52, 24], or to reconstruct the unlearned data [56, 57].
@@ V-A Audit of Privacy Leakage
V-B Assessment of Model Robustness
Similarly, malicious unlearning can assess the robustness of the unlearned model. The foundational principle of this assessment approach is akin to that used in adversarial attacks, as detailed in works such as [102, 103, 104, 105]. For example, [23] describes a method of malicious unlearning via over-unlearning, where perturbed data are submitted in the unlearning request to alter the classifier’s decision boundary. This strategy of pushing the decision boundary allows for evaluating the impact of unlearned data points near the boundary on model performance, as well as the effect of data points that challenge the model’s prediction capabilities. Another example occurs in the context of LLMs[57], where attackers insert adversarial embeddings into the prompts’ embeddings, aiming to extract supposedly unlearned knowledge from the model. Such attacks provide a means to gauge the model’s resilience against malicious input embeddings. On a different note, [106] and [107] treat data removal as an adversarial perturbation to the overall training dataset. In this case, the concept of randomized smoothing [108] can be adopted to measure the model robustness against unlearned data [106] or unlearned users in a federated setting [107], hence establishing a certified budget for data removals, within which the model retains its robustness.
V-C Proof of Unlearning
@@
VI Challenges and Promising Directions
@@
VI-A Defenses against Malicious Unlearning
Note that one of the main threats to machine unlearning systems is the negative impact on unlearned models caused by malicious unlearning (see Section III-B for more details). This threat manifests as either degraded model performance or the introduction of backdoors, involving the submission of malicious unlearning requests. While direct malicious unlearning requests can often be easily detected by checking if the requested data for removal is part of the training dataset [23], identifying such requests during preconditioned unlearning attacks presents a nontrivial challenge. This is because the data requested for removal in these attacks are actually part of the training dataset and usually are mitigation data without adversarial features. Therefore, developing robust detection mechanisms for identifying these subtle and sophisticated unlearning requests becomes essential in safeguarding the integrity of machine unlearning systems.
VI-B Federated Unlearning
@@ VI-C Privacy Preservation
VI-D Unlearning for Large Models
@@
The inefficiency of the unlearning for large models, which calls for innovative solutions to improve its effectiveness [132].

### v4

@@ VII-B Value Alignment
In these efforts, unlearning is applied in various innovative ways to align AI operations with ethical and legal standards. For instance, attribute unlearning [106] is used to remove sensitive attributes for privacy compliance. Similarly, unlearning with corresponding detection methods are applied to LLMs to remove behaviors and content that are deemed inappropriate or should be avoided, effectively eliminating illegal, poor quality, copyrighted, or harmful content [34, 107, 108, 109, 110, 111, 112, 113]. In a separate vein, the work outlined by Isonuma et al. [114] introduces a distinctive application of unlearning principles aimed at identifying the training datasets that have the most profound influence on generating harmful content. Through a process of systematically unlearning each dataset via gradient ascent, the repercussions on the model’s propensity to generate harmful content can be meticulously evaluated post-unlearning.
VIII Evaluating Unlearning through Attacks
@@
VIII-A Audit of Privacy Leakage
As outlined in Section VI-A, machine unlearning systems are susceptible to threats of information leakage. Attackers can exploit discrepancies between the trained and unlearned models, or inherent vulnerabilities within the unlearning system, to initiate inference attacks [55, 66, 67, 51], or to reconstruct the unlearned data [71, 72].
@@ VIII-A Audit of Privacy Leakage
VIII-B Assessment of Model Robustness
Similarly, malicious unlearning can assess the robustness of the unlearned model. The foundational principle of this assessment approach is akin to that used in adversarial attacks, as detailed in works such as [118, 119, 120, 121]. For example, [50] describes a method of malicious unlearning via over-unlearning, where perturbed data are submitted in the unlearning request to alter the classifier’s decision boundary. This strategy of pushing the decision boundary allows for evaluating the impact of unlearned data points near the boundary on model performance, as well as the effect of data points that challenge the model’s prediction capabilities. Another example occurs in the context of LLMs[72], where attackers insert adversarial embeddings into the prompts’ embeddings, aiming to extract supposedly unlearned knowledge from the model. Such attacks provide a means to gauge the model’s resilience against malicious input embeddings. On a different note, [122] and [123] treat data removal as an adversarial perturbation to the overall training dataset. In this case, the concept of randomized smoothing [124] can be adopted to measure the model robustness against unlearned data [122] or unlearned users in a federated setting [123], hence establishing a certified budget for data removals but retains the model robustness.
VIII-C Proof of Unlearning
@@
IX Summary and Discussion
So far, we have reviewed the related works on threats in MU systems, along with an analysis of various attack types, defense mechanisms, and their interactions in different roles. To clarify these relationships, as illustrated in Figure 5, this section offers a summary and discussion.
Machine unlearning. MU presents both threats and opportunities for defense: (i) the process of machine unlearning and the handling of unlearning requests introduce threats such as information leakage and the potential for malicious unlearning attacks, and (ii) unlearning can be used as a defensive strategy for model recovery and remove undesirable knowledge for value alignment.
Threats. Threats arise from the unlearning process and the workflow of MU systems, creating opportunities for attacks. Therefore, corresponding defensive strategies must be explored and developed.
Attacks. Attacks serve two key purposes: (i) to explore potential threats in MU systems, and (ii) to evaluate the performance of unlearning systems, including aspects such as privacy leakage, robustness, and effectiveness.
Defenses. As mentioned earlier, (i) MU can serve as a defense in certain scenarios, and (ii) defenses against specific threats in MU systems still need to be thoroughly explored.
X Challenges and Promising Directions
@@
X-A Defenses against Malicious Unlearning
Note that one of the main threats to machine unlearning systems is the negative impact on unlearned models caused by malicious unlearning (see Section VI-B for more details). This threat manifests as either degraded model performance or the introduction of backdoors, involving the submission of malicious unlearning requests. While direct malicious unlearning requests can often be easily detected by checking if the requested data for removal is part of the training dataset [50], identifying such requests during preconditioned unlearning attacks presents a nontrivial challenge. This is because the data requested for removal in these attacks are actually part of the training dataset and usually are mitigation data without adversarial features. Therefore, developing robust detection mechanisms for identifying these subtle and sophisticated unlearning requests becomes essential in safeguarding the integrity of machine unlearning systems.
X-B Federated Unlearning
@@ X-C Privacy Preservation
X-D Unlearning for Large Models
@@
The inefficiency of the unlearning for large models, which necessitates innovative solutions to improve its effectiveness [35].